For the following items, find the sinking fund payment

5.) Cassius landscaping buys new lawn equipment every three years. It is estimated that 250000 will be needed for the next purchase. The company sets up a sinking fund to save for this obligation.a) what equal payments must be deposited every 6 months, if interest is 8% compounded semi annually ? B) what is the total amount of interest earned by the sinking fund?


1
Expert's answer
2021-05-21T00:13:02-0400

(a) The sinking fund is for 6 period ( 3 years"\times" 2 period per year) at 4% per period (8%/2 period per year).


 From table 12-1, 6 period, 4% per period gives a future value table factor of 6.63298.


sinking fund payment "= \dfrac{\text{ Future value of sinking fund}}{\text{ Future value table factor}}"

       

"=\dfrac{25000}{6.63298}=3769.04"


(b) The total payment have to pay in 6 period

Total payment ="3769\times 6"

           =$22614.24

Therefore, interest earn = $(25000-22614.24)

             =$2385.7